# Spin Theory Based on the Extended Least Action Principle and Information Metrics: Quantization, Entanglement, and Bell Test With Time Delay

### v2

@@ Abstract
A theory of electron spin is developed here based on the extended least action principle and assumptions of intrinsic angular momentum of an electron with random orientations. By incorporating appropriate relative entropy for the random orientations of intrinsic angular momentum in the extended least action principle, the theory recovers the quantum formulation of electron spin. The two-level quantization of spin measurement is a natural mathematical consequence. The formulation of measurement probability when a second Stern-Gerlach apparatus is rotated relative to the first Stern-Gerlach apparatus, and the Schrödinger-Pauli equation, are also derived successfully. Furthermore, we provide an intuitive physical model and formulation to explain the entanglement phenomenon between two electron spins. In this model, spin entanglement is the consequence of correlation between the random orientations of the intrinsic angular momenta of the two electrons. Since the orientation is an intrinsic local property of electron, the correlation of orientations can be preserved even when the two electrons are remotely separated. Such a correlation can be manifested without causal effect. Owing to this orientation correlation, the Bell-CHSH inequality is shown to be violated in a Bell test. A potential experiment is proposed to test the validity of the present theory. In a typical Bell test that confirms the violation of Bell-CHSH inequality, the present theory suggests that by adding a sufficiently large time delay before Bob’s measurement, the Bell-CHSH inequality can become non-violated. Limitations of the theory are also discussed to inspire further investigations on a more rigorous physical model of electron spin.
@@ I Introduction
In quantum mechanics, the spin of an electron is one of the most important physical observables to demonstrate the essences of quantum theory. These essential features include, for instance, 1.) the quantization of measurement outcome - measuring the spin of an electron always obtains two possible values, spin-up or spin-down, along the direction of measuring magnetic field; 2.) the need of multi-components wave functions in the Schrödinger-Pauli equation to describe the dynamics of an electron with spin; 3.) quantum entanglement - the singlet state of an electron pair is often used as example to illustrate the conceptual challenges such as in the Bohm version EPR thought experiment EPR ; Bohm1957 , or to verify the violation of Bell inequality Bell . However, these important quantum features are either not derived from first principles or still have conceptual difficulty. Quantization of spin measurement outcome is normally introduced as postulate in standard quantum mechanics; Whether the wave function is epistemological or associated with physical reality is still under debate; The violation of Bell inequality confirms that spin correlation can be non-local in the sense that the correlation is inseparable even if the two electrons are space-like separated. Without an intuitive physical model, such inseparability is not possible to be comprehended in classical terms and still puzzling the physics community Brunner . A more intuitive physical model and a clearer theory of electron spin to better explain these challenges are still much desirable in modern quantum physics.
@@
Historically, a number of physical models for electron spin were proposed Tomonaga ; Sebens . Here we will briefly review two prominent ones. The first model is to consider the electron as a rotating rigid body with uniform distributed charge, and has been adopted in stochastic mechanism Dankel1977 ; Fairs1982 ; Beyer . This model can calculate the relation between angular momentum and magnetic moment of an electron. By applying the theory of stochastic mechanics, the electron spin is treated as the averaged of hidden angular momentum variables with random orientations. Formulations equivalent to quantum mechanics can be recovered. However, one of the problems for this model is that when the radius of the rigid body is sufficiently small, the edge is moving faster than the speed of light Tomonaga . In addition, the velocity fields in stochastic mechanics retain the non-local characteristics, a drawback that we attempt to avoid. Another model is the so-called “Zitterbewegung” model where the origin of spin is proposed as a result of helical motion of a ”light-like” particle Wey1947 ; Barut1984 ; Niehaus16 ; Niehaus22 . The particle is moving automatically in a circle with the speed of light. There is also a simpler model that considers an electron as a point particle with intrinsic magnetic moment and intrinsic angular momentum. Such a model offers no explanation on how the intrinsic angular moment or intrinsic magnetic moment originate.
@@
Recent interests in searching for foundational principles of quantum theory from the information perspective Rovelli:1995fv ; zeilinger1999foundational ; Brukner:ys ; Brukner:1999qf ; Brukner:2002kx ; Fuchs2002 ; brukner2009information ; Brukner:vn ; spekkens2007evidence ; Spekkens:2014fk ; Paterek:2010fk ; gornitz2003introduction ; lyre1995quantum ; Hardy:2001jk ; masanes2011derivation ; Mueller:2012ai ; Masanes:2012uq ; chiribella2011informational ; Mueller:2012pc ; Hardy:2013fk ; kochen2013reconstruction ; 2008arXiv0805.2770G ; Hall2013 ; Hoehn:2014uua ; Hoehn:2015 ; Stuckey ; Mehrafarin2005 ; Caticha2011 ; Frieden ; Reginatto can offer new perspectives for the investigations of spin theory. One of the motivations is to reformulate quantum mechanics with information theoretic principles such that some of the conceptual challenges can be resolved. Inspired by this program, an extended least action principle has been proposed Yang2023 . The principle extends the classical least action principle by incorporating information metrics about vacuum fluctuations. Based on this principle, the non-relativistic quantum mechanics Yang2023 and quantum scalar field theory Yang2024 have been recovered, demonstrating the general applicability of the theoretical framework.
@@
The goal of this paper is to apply the extended least action principle to develop the theory for electron spin with the goal of resolving some of the challenges mentioned earlier. This is achieved by introducing an additional assumption of intrinsic angular momentum with random orientations, and choosing proper relative entropy for the random orientations. We will show the results are indeed quite fruitful. Discreteness of measurement outcome is a natural mathematical result when the order of relative entropy approaches a limit. Formulation equivalent to quantum mechanics is obtained when the direction of magnetic field of a second Stern-Gerlach apparatus is rotated with an angle from the direction of magnetic field of the first Stern-Gerlach apparatus. Recursively applying the extended least action principle, we can derive the Schrödinger-Pauli equation.
@@
With respect to spin entanglement between two electron spins, we will show that the root cause of entanglement is the correlation between the random variables, the orientations of the intrinsic angular momenta, of the two electrons. The correlation can be established during preparation. Crucially, since the orientation of the angular momentum is an intrinsic local property of the electron, the correlation can be maintained even though the two electrons are remotely separated and without interaction. Mathematically, we give an equivalent formulation of the four Bell states using probability density function instead of wave function. Based on the probability density function, Bell-CHSH inequality is proved to be violated in our theory due to the same root cause of entanglement. Since the correlation of the orientations of the angular momenta is preserved even when the two electrons are separated, the joint probability for the measurement outcome cannot be factorized into a product of two individual terms - a requirement for the Bell inequality to hold.
@@
The theory gives certain results that slightly different from standard quantum mechanics. Two potential experiments are proposed to test the difference and the validity of the theory. In the first experiment, the Stern-Gerlach apparatus is configured with a sufficiently weak gradient of the inhomogeneous magnetic field along the $z$-axis; We expect the electron detector screen will exhibit continuous distribution of displacements along the $z$-axis around two areas, instead of only two discrete lines. The second experiment is to modify a typical Bell test experiment that has already confirmed the violation of Bell-CSHS inequality. Here, we purposely let Alice and Bob perform their measurements at different times on each entangled pair of electrons. While Alice performs her measurement at time $t_{a}$, Bob purposely delays his measurement at time $t_{b}=t_{a}+\Delta t$. The present theory suggests that when $\Delta t$ is chosen properly, the Bell-CSHS inequality will not be violated.
@@
The rest of the paper is organized as follows. In Section II, we review the extended least action principle and its underlying assumptions. Detailed calculation showing how the basic quantum theory is derived from the principle is provided in Appendix A. Section III is devoted to developing the spin model, where the extended least action principle is incorporated with Tsallis relative entropy, resulting in a probability density function for the random orientation. The probability density function becomes two Dirac delta functions when the order of relative entropy approaches a limit, corresponding to the quantization of measurement outcomes. The theory for rotation of the spin is developed in this section as well. In Section IV, the Schrödinger-Pauli equation is derived by combining the effects of random translational fluctuations and random rotational fluctuations when applying the extended least action principle. Section V develops the theory for spin entanglement based on the correlations between the random orientations of intrinsic angular momenta, which allows us to prove that Bell inequality is violated with this spin theory. To test the validity of the theory, two potential experiments are proposed in Section VI.
@@
There are clear limitations in the present theory, particularly the assumptions on the detailed physical model of electron spin, and the assumption that the order of relative entropy increases when the electron travels along an inhomogeneous magnetic field. A more rigorous theory must be developed to further justify these assumptions and beyond the scope of the present theory. The limitations of the present theory are discussed in Section VII.
@@ II The Extended Least Action Principle
The theoretical framework in this paper is built on the extended least action principle proposed in Ref. Yang2023 . We will briefly review the principle before introducing additional assumptions in order to develop the spin theory. In Yang2023 , the least action principle in classical mechanics is extended to derive quantum formulation by factoring in the following two assumptions.
@@
The first assumption is generally accepted in mainstream quantum mechanics, which is responsible for the intrinsic randomness of the dynamics of a quantum object. Even though we do not know the physical details of the vacuum fluctuation, the crucial assumption here is the locality of the vacuum fluctuation. This implies that for a composite system, the fluctuation of each subsystem is independent of each other.
@@
The implications of the second assumption need more elaborations. Historically the Planck constant was first introduced to show that energy of radiation from a black body is discrete. One can consider the discrete energy unit as the smallest unit to be distinguished, or detected, in the black body radiation phenomenon. In general, it is understood that Planck constant is associated with the discreteness of certain observables in quantum mechanics. Here, we instead interpret the Planck constant from an information measure point of view. Assumption 2 states that there is a lower limit to the amount of action that the physical system needs to exhibit in order to be observable or distinguishable in potential observation, and such a unit of action is determined by the Planck constant.
@@
Making use of this understanding of the Planck constant conversely provides us a new way to calculate the additional action due to vacuum fluctuations. That is, even though we do not know the physical details of vacuum fluctuations, the vacuum fluctuations manifest themselves via a discrete action unit determined by the Planck constant as an observable information unit. If we are able to define an information metric that quantifies the amount of observable information manifested by vacuum fluctuations, we can then multiply the metric with the Planck constant to obtain the action associated with vacuum fluctuations. Then, the challenge to calculate the additional action due to vacuum fluctuation is converted to define a proper new information metric $I_{f}$, which measures the additional distinguishable, hence observable, information exhibited due to vacuum fluctuations. Even though we do not know the physical details of vacuum fluctuations (except that as Assumption 1 states, these vacuum fluctuations are completely random and local), the problem becomes less challenged since there are information-theoretic tools available.
@@
The first step is to assign a transition probability distribution due to vacuum fluctuation for an infinitesimal time step at each position along the classical trajectory. The distinguishability of vacuum fluctuation then can be defined as the information distance between the transition probability distribution and a uniform probability distribution. Uniform probability distribution is chosen here as reference to reflect the complete randomness of vacuum fluctuations. In information theory, the common information metric to measure the information distance between two probability distributions is relative entropy. Relative entropy is more fundamental to Shannon entropy since the latter is just a special case of relative entropy when the reference probability distribution is a uniform distribution. But there is a more important reason to use relative entropy. As shown in later sections, when we consider the dynamics of the system for an accumulated time period, we assume the initial position is unknown but is given by a probability distribution. This probability distribution can be defined along the position of classical trajectory without vacuum fluctuations, or with vacuum fluctuations. The information distance between the two probability distributions gives the additional distinguishability due to vacuum fluctuations. It is again measured by a relative entropy. Thus, relative entropy is a powerful tool allowing us to extract meaningful information about the dynamic effects of vacuum fluctuations. Concrete form of $I_{f}$ will be defined later as a functional of relative entropy that measures the information distances of different probability distributions caused by vacuum fluctuations. Thus, the total action from classical path and vacuum fluctuation is
@@
where $S_{c}$ is the classical action. Non-relativistic quantum theory can be derived through a variation approach to minimize such a functional quantity, $\delta S_{t}=0$. When $\hbar\to 0$, $S_{t}=S_{c}$. Minimizing $S_{t}$ is then equivalent to minimizing $S_{c}$, resulting in Newton’s laws in classical mechanics. However, in quantum mechanics, $\hbar\neq 0$, the contribution from $I_{f}$ must be included when minimizing the total action. We can see $I_{f}$ is where the quantum behavior of a system comes from. These ideas can be condensed as
@@
Non-relativistic quantum formulation can be derived based on the extended least action principle if we only consider the translational component of vacuum fluctuations Yang2023 . A brief description of the derivation is given in Appendix A for self reference. It has also been shown that the quantum scalar field theory can be formulated based on the principle Yang2024 , which further demonstrates its general applicability. We will next apply the extended least action principle to derive the quantum theory of electron spin by considering not only translational, but rotational components of vacuum fluctuations. It turns out that the results are more fruitful than the results already obtained in Yang2023 ; Yang2024 .

### v3

# Quantum Theory of Electron Spin Based on the Extended Least Action Principle and Information Metrics of Spin Orientations

@@ Abstract
Quantum theory of electron spin is developed here based on the extended least action principle and assumptions of intrinsic angular momentum of an electron with random orientations. The novelty of the formulation is the introduction of relative entropy for the random orientations of intrinsic angular momentum when extremizing the total actions. Applying recursively this extended least action principle, we show that the quantization of electron spin is a mathematical consequence when the order of relative entropy approaches a limit. In addition, the formulation of the measurement probability when a second Stern-Gerlach apparatus is rotated relative to the first Stern-Gerlach apparatus, and the Schrödinger-Pauli equation, are recovered successfully. Furthermore, the principle allows us to provide an intuitive physical model and formulation to explain the entanglement phenomenon between two electron spins. In this model, spin entanglement is the consequence of the correlation between the random orientations of the intrinsic angular momenta of the two electrons. Since spin orientation is an intrinsic local property of the electron, the correlation of spin orientations can be preserved and manifested even when the two electrons are remotely separated. The entanglement of a spin singlet state is represented by two joint probability density functions that reflect the orientation correlation. Using these joint probability density functions, we prove that the Bell-CHSH inequality is violated in a Bell test. To test the validity of the spin-entanglement model, we propose a Bell test experiment with time delay. Such an experiment starts with a typical Bell test that confirms the violation of the Bell-CHSH inequality but adds an extra step that Bob’s measurement is delayed with a period of time after Alice’s measurement. The present theory suggests that the Bell-CHSH inequality becomes non-violated if the time delay is sufficiently large.
@@ I Introduction
In quantum mechanics, the spin of an electron is one of the most important physical observables to demonstrate the essence of quantum theory. For instance, measurement of the spin of an electron always obtains two possible values, spin-up or spin-down, along the direction of the measuring magnetic field; To describe the dynamics of an electron with spin, multi-components wave functions in the Schrödinger-Pauli equation are needed; In investigating the quantum entanglement phenomenon, the singlet state of an electron pair is often used as example to illustrate conceptual challenges such as in the Bohm version EPR thought experiment EPR ; Bohm1957 , or to verify the violation of Bell inequality Bell . However, these important quantum features are either not derived from first principles or still have conceptual difficulties. Quantization of spin measurement outcome is normally introduced as a postulate in standard quantum mechanics. Whether the wave function is epistemological or associated with physical reality is still under debate. The violation of Bell inequality confirms that spin correlation can be non-local in the sense that the correlation is inseparable even if the two electrons are space-like separated. Without an intuitive physical model, this inseparability is not possible to comprehend in classical terms and still puzzles the physics community Brunner . A more intuitive physical model and a clearer theory of electron spin to better explain these challenges are still much desirable in modern quantum physics.
@@
Historically, a number of physical models for electron spin have been proposed Tomonaga ; Sebens . Here we will briefly review two prominent ones. The first model considers the electron as a rotating rigid body with a uniform distributed charge, and has been adopted in a stochastic mechanism Dankel1977 ; Fairs1982 ; Beyer . This model can derive the relation between the angular momentum and the magnetic moment of an electron. By applying the theory of stochastic mechanics, the electron spin is treated as the averaged of hidden angular momentum variables with random orientations. Formulations equivalent to quantum mechanics can be recovered. However, one of the problems of this model is that when the radius of the rigid body is small enough, the edge moves faster than the speed of light Tomonaga . In addition, the velocity fields in stochastic mechanics retain the non-local characteristics, a drawback that should be avoided. Another model is the so-called “Zitterbewegung” model where the origin of spin is proposed as a result of the helical motion of a ”light-like” particle Wey1947 ; Barut1984 ; Niehaus16 ; Niehaus22 . The particle is moving automatically in a circle with the speed of light. There is also a simpler model that considers an electron as a point particle with intrinsic magnetic moment and intrinsic angular momentum. Such a model does not offer an explanation as to how the intrinsic angular moment or the intrinsic magnetic moment originates.
@@
Recent interest in searching for foundational principles of quantum theory from the information perspective Rovelli:1995fv ; zeilinger1999foundational ; Brukner:ys ; Brukner:1999qf ; Fuchs2002 ; brukner2009information ; spekkens2007evidence ; Spekkens:2014fk ; Paterek:2010fk ; gornitz2003introduction ; lyre1995quantum ; Hardy:2001jk ; masanes2011derivation ; Mueller:2012ai ; Masanes:2012uq ; chiribella2011informational ; Mueller:2012pc ; Hardy:2013fk ; kochen2013reconstruction ; 2008arXiv0805.2770G ; Hall2013 ; Hoehn:2014uua ; Hoehn:2015 ; Caticha2011 ; Caticha2019 ; Frieden ; Reginatto can offer new perspectives for the investigation of spin theory. The motivation here is to reformulate quantum mechanics with information-theoretic principles such that some of the conceptual challenges can be resolved. Inspired by this program, an extended least action principle has been proposed Yang2023 . The principle extends the classical least action principle by incorporating information metrics on vacuum fluctuations. Based on this principle, non-relativistic quantum mechanics Yang2023 and quantum scalar field theory Yang2024 have been recovered, demonstrating the general applicability of the theoretical framework.
@@
The goal of this paper is to apply the extended least action principle to develop the theory for electron spin with the goal of providing new insights into some of the challenges mentioned earlier. This is achieved by introducing an additional assumption of intrinsic angular momentum with random orientations and choosing appropriate relative entropy for the random orientations. We will show that the results are indeed quite fruitful. Discreteness of measurement result of electron spin is a mathematical result when the order of relative entropy approaches a limit. The formulation also gives equivalent results as standard quantum mechanics when the direction of magnetic field of a second Stern-Gerlach apparatus is rotated with an angle from the direction of magnetic field of the first Stern-Gerlach apparatus. Recursively applying the extended least action principle, we can derive the Schrödinger-Pauli equation.
@@
With respect to spin entanglement between two electron spins, we will demonstrate that the origin of entanglement is attributed to the correlation between the random variables, the orientations of the intrinsic angular momenta, of the two electrons. The correlation can be established during preparation. The correlation can be preserved even though the two electrons are remotely separated and without interaction since the orientation of the angular momentum is an intrinsic local property of the electron. Mathematically, we give an equivalent formulation of the four Bell states using the probability density function of the angular momentum orientation, instead of the wave function in standard quantum mechanics. Based on the probability density function of a spin singlet state, the Bell-CHSH inequality is proved to be violated due to the same root cause of entanglement - the correlation between the random orientations of the intrinsic angular momenta. Specifically, since this correlation is preserved even when the two electrons are separated, the joint probability for the measurement outcome cannot be factorized into a product of two individual terms, a requirement for the Bell inequality to hold.
@@
Two potential experiments are proposed to test the validity of the theory and its difference from standard quantum mechanics. In the first experiment, the Stern-Gerlach apparatus is configured with a sufficiently weak gradient of the inhomogeneous magnetic field along the $z$-axis; We expect the electron detector screen to exhibit a continuous distribution of displacements along the $z$-axis around two areas, instead of only two discrete lines. The second experiment is to modify a typical Bell test experiment that has already confirmed the violation of the Bell-CSHS inequality. Here, we purposely let Alice and Bob perform their measurements at different times on each entangled pair of electrons. While Alice performs her measurement at time $t_{a}$, Bob intentionally delays his measurement at time $t_{b}=t_{a}+\Delta t$. The present theory suggests that when $\Delta t$ is chosen properly, the Bell-CSHS inequality will not be violated.
@@
The paper is organized as follows. In Section II, we review the extended least action principle and its underlying assumptions. Detailed calculations showing how the basic quantum theory is derived from the principle are provided in Appendix A. Section III is devoted to developing the spin model, where the extended least action principle is incorporated with the relative entropy for the spin random orientations, resulting in a probability density function for the spin orientation. The probability density function becomes two Dirac delta functions when the order of relative entropy approaches a limit, corresponding to the quantization of measurement outcomes. The theory for rotation of the spin is developed in this section as well. In Section IV, the Schrödinger-Pauli equation is derived by combining the effects of random translational fluctuations and random rotational fluctuations when applying the extended least action principle. Section V develops the theory for spin entanglement based on the correlations between the random orientations of intrinsic angular momenta, which allows us to prove that the Bell inequality is violated for a singlet state. To test the validity of the theory, two potential experiments are proposed in Section VI.
@@
There are clear limitations in the present theory, as discussed in Section VII, particularly the assumption that the order of relative entropy increases when the electron travels along an inhomogeneous magnetic field. We give some intuitive explanations in Section VII, but a more rigorous theory must be developed to further justify these assumptions.
@@ II The Extended Least Action Principle
The theoretical framework in this paper is developed based on the extended least action principle proposed in Yang2023 . We will briefly review the principle before introducing additional assumptions in order to develop the spin theory. In Yang2023 , the least action principle in classical mechanics is extended to derive the quantum formulation by factoring in the following two assumptions.
@@
The first assumption is generally accepted in mainstream quantum mechanics and is responsible for the intrinsic randomness of the dynamics of a quantum object. Although we do not know the physical details of the vacuum fluctuation, the crucial assumption here is the locality of the vacuum fluctuation. This implies that for a composite system, the fluctuation of each subsystem is independent of each other.
@@
The implications of the second assumption need more elaborations. Historically, the Planck constant was first introduced to show that the energy of radiation from a black body is discrete. One can consider the discrete energy unit as the smallest unit to be distinguished, or detected, in the black-body radiation phenomenon. In general, it is understood that Planck constant is associated with the discreteness of certain observables in quantum mechanics. Here, we instead interpret the Planck constant from an information measure point of view. Assumption 2 states that there is a lower limit to the amount of action that the physical system needs to exhibit in order to be observable or distinguishable in potential observation, and such a unit of action is determined by the Planck constant.
@@
Making use of this understanding of the Planck constant conversely provides us a new way to calculate the additional action due to vacuum fluctuations. That is, even though we do not know the physical details of vacuum fluctuations, the vacuum fluctuations manifest themselves via a discrete action unit determined by the Planck constant as an observable information unit. If we can define an information metric that quantifies the amount of observable information manifested by vacuum fluctuations, we can then multiply the metric by the Planck constant to obtain the action associated with vacuum fluctuations. Then, the challenge of calculating the additional action due to the vacuum fluctuation is converted to define a proper new information metric $I_{f}$, which measures the additional distinguishable, and hence observable, information exhibited due to vacuum fluctuations. The problem of defining an appropriate information metrics becomes less challenging since there are information-theoretic tools available.
@@
The first step is to assign a transition probability distribution due to vacuum fluctuation for an infinitesimal time step at each position along the classical trajectory. The distinguishability of vacuum fluctuation can then be defined as the information distance between the transition probability distribution and a uniform probability distribution. The uniform probability distribution is chosen here as a reference to reflect the complete randomness of vacuum fluctuations. In information theory, the common information metric for measuring the distance between two probability distributions is relative entropy. Relative entropy is more fundamental than Shannon entropy since the latter is just a special case of relative entropy when the reference probability distribution is a uniform distribution. However, there is a more important reason to use relative entropy. As shown in later sections, when we consider the dynamics of the system for an accumulated time period, we assume that the initial position is unknown but given by a probability distribution. This probability distribution can be defined along the position of a classical trajectory without vacuum fluctuations or with vacuum fluctuations. The information distance between the two probability distributions gives additional distinguishability because of vacuum fluctuations. It is again measured by a relative entropy. Thus, relative entropy is a powerful tool allowing us to extract meaningful information about the dynamic effects of vacuum fluctuations. The concrete form of $I_{f}$ will be defined later as a functional of relative entropy that measures the information distances of different probability distributions caused by vacuum fluctuations. Thus, the total action from classical path and vacuum fluctuation is
@@
where $S_{c}$ is the classical action. Non-relativistic quantum theory can be derived through a variation approach to minimize such a functional quantity, $\delta S_{t}=0$. When $\hbar\to 0$, $S_{t}=S_{c}$. Minimizing $S_{t}$ is then equivalent to minimizing $S_{c}$, resulting in Newton’s laws in classical mechanics. However, in quantum mechanics, $\hbar\neq 0$, the contribution of $I_{f}$ must be included when minimizing the total action. We can see that $I_{f}$ is where the quantum behavior of a system comes from. These ideas can be condensed as follows.
@@
Non-relativistic quantum formulation can be derived based on the extended least action principle if we only consider the translational component of vacuum fluctuations Yang2023 . A brief description of the derivation is given in Appendix A for self-reference. It has also been shown that the quantum scalar field theory can be formulated based on the principle Yang2024 , which further demonstrates its general applicability. In this paper, we will apply the extended least action principle to derive the quantum theory of electron spin by considering not only translational, but also rotational components of vacuum fluctuations.